Melissa is 24 years younger than Joyce. In 2 years Joyce will be 3 times as old as Melissa will be then. How old are they now?
Answer by stanbon(75887)

Melissa is 24 years younger than Joyce. In 2 years Joyce will be 3 times as old as Melissa will be then. How old are they now?
----
Equations:
m = j-24
(j+2)= 3(m+2)
-------------------
Substitute for "m" and solve for "j":
j+2 = 3(j-24+2)
j+2 = 3(j-22)
j+2 = 3j - 66
2j = 68
j = 34 (Joyce's age now)
-------------------------
Since m = j-24, m = 34-24 = 10 yrs old (Melissa's age now)
===========================================================
Cheers,
Stan H.
============================